OpenStudy (anonymous):

What is the length of the segment with endpoints A(1,7) and B(-3, -1)?

TheSmartOne (thesmartone):

Hi there! :)

TheSmartOne (thesmartone):

Do you know the distance formula? (:

OpenStudy (anonymous):

I know it but I just forget it @TheSmartOne

TheSmartOne (thesmartone):

\(\LARGE Distance = \sqrt{(x_2 - x_1)^2+(y_2 - y_1)^2}\)

OpenStudy (alexandervonhumboldt2):

substitute your points and solve

TheSmartOne (thesmartone):

You think you can do the rest from here on? :)

OpenStudy (anonymous):

Thank you.. I can figure it out now @TheSmartOne

TheSmartOne (thesmartone):

No problem, :)

OpenStudy (anonymous):

Would you happen to know what the midpoint formula is @TheSmartOne

OpenStudy (anonymous):

Square root of 80

TheSmartOne (thesmartone):

@kenna_098 Surprising that I didn't get a notification about your tag, mhmm. The midpoint formula is: \(\sf\Huge (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\) that would be the midpoint of the two points (x1,y1) and (x2,y2)